OpenStudy (anonymous):

Hello there having some trouble with some integral quesitons. Help appreacited and medal of course. Will post below

OpenStudy (anonymous):

Find the derivative of \[F(x)=\int\limits_{\ln(x)}^{\ln(x+1)}2\sinh(t)dt\]

OpenStudy (anonymous):

So i try doing the question using the fact that sinh=\[(e^x-e^-x)/2, \] so the 2 will cancel, then I try using the fundamental therom and chain rule but it still isnt working.

OpenStudy (anonymous):

Any insight would be great, thanks

OpenStudy (anonymous):

i would say \(\sinh(\ln(x+1))\times\frac{1}{x+1}-\sinh(\ln(x))\times \frac{1}{x}\) although there might be a slicker way of writing it

OpenStudy (anonymous):

damn i forgot the \(2\)

OpenStudy (anonymous):

oh your method is much better

OpenStudy (anonymous):

Hey yea well I know the answer is given as 2x+1/(x)^2(x+1)^2 but I cant seem to get it into that form

OpenStudy (anonymous):

lets try it

OpenStudy (anonymous):

\[\frac{e^{\ln(x+1)}-e^{-\ln(-x-1)}}{x+1}\] for the first part

OpenStudy (anonymous):

that gives \[\frac{x+1-\frac{1}{x+1}}{x+1}\]

OpenStudy (anonymous):

damn i made a typo above, but the second part is right

OpenStudy (anonymous):

second part will be \[-\frac{x+\frac{1}{x}}{x}\]

OpenStudy (anonymous):

then i guess a raft of algebra

OpenStudy (anonymous):

first line should have been \[\frac{e^{\ln(x+1)}-e^{-\ln(x+1)}}{x+1}\]

OpenStudy (31356):

Correct!

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

no way i can do all the algebra on the fly, but did you get as far as \[\frac{x+1-\frac{1}{x+1}}{x+1}-\frac{x+\frac{1}{x}}{x}\]?

OpenStudy (anonymous):

Yea thanks for the help. I guess that answer would be accepted anyways

OpenStudy (anonymous):

algebra from here on in remove the complex fractions, then subtract